### v2

# CoSQA+: Enhancing Code Search Dataset with Matching Code

Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets are problematic: either using unrealistic queries, or with mismatched codes, and typically using one-to-one query-code pairing, which fails to reflect the reality that a query might have multiple valid code matches. This paper introduces CoSQA+, pairing high-quality queries (reused from CoSQA) with multiple suitable codes. We collect code candidates from diverse sources and form candidate pairs by pairing queries with these codes. Utilizing the power of large language models (LLMs), we automate pair annotation, filtering, and code generation for queries without suitable matches. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. Furthermore, we propose a new metric Mean Multi-choice Reciprocal Rank (MMRR), to assess one-to-N code search performance. We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
Introduction
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development  (Li et al. 2013; Sim et al. 2011; Stolee, Elbaum, and Dobos 2014). Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems  (Yan et al. 2020). Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. For the practical usability of neural code search, real user queries are crucial to a high-quality dataset  (Sun et al. 2022). The key challenge is how to pair real-world queries with codes for a high-quality code search benchmark.
However, there are some common challenges in existing code search datasets. Datasets like CodeSearchNet (Husain et al. 2019) contain general-purpose queries from GitHub, but these queries are either documentation strings or comments rather than natural language questions. Models trained on these unrealistic queries often fail to generalize to real-world use cases, leading to poor performance in practical applications (Liu et al. 2009; Xie et al. 2023). CoSQA contains realistic queries from the Bing search engine, but we found that around 51% of queries are paired with mismatching code, which adversely affects the training of code search models. When code snippets do not correspond accurately to the query descriptions, this leads to excessive noise in the dataset, complicating the model’s ability to retrieve relevant code (Li, Zhou, and Shen 2024; Sun et al. 2022). Additionally, existing code search datasets typically pair each query with a single code snippet, which limits the evaluation of queries that could match multiple code snippets. Finally, building query-code datasets involves laboriously matching and labeling real search queries with accurate code. Despite employing 100 human annotators to cover 20,604 queries, CoSQA only managed to generate 6,267 unique code snippets, with half of the queries failing to match the correct code.
To address these problems, we develop a new benchmark CoSQA+, pairing high-quality queries with multiple high-quality codes. The queries come from CoSQA, and the codes include snippets, blocks and functions from the filtered StaQC and CodeSearchNet datasets. The candidate pairs are formed by pairing the queries with the codes through multiple models. Utilizing powerful Claude 3 Sonnet, we automatically label the candidate pairs with CoT prompts. Given the limited size of our code pool, some queries fail to be matched with suitable code snippets. We use GPT-4o to generate exact matching codes for these queries and manually verify their accuracy. Furthermore, a new metric, MMRR, is proposed to evaluate models’ performance on one-to-N code search.
We conduct four experiments to assess the quality of CoSQA+. (1) In a quality comparison between CoSQA+ and CoSQA, for the 1000 queries, 62.9% of the paired codes from CoSQA+ are selected as better. (2) When CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings are fine-tuned on CoSQA+, they all perform better on the CSN Python dataset compared to being fine-tuned on CoSQA. (3) Automated Claude 3 Sonnet annotation yields performance close to human levels, with a Krippendorff’s Alpha of 0.628 and an accuracy of 84% in exact match conditions. (4) The MMRR metric proves to be highly reliable and stable for evaluating the effectiveness of multi-choice code search on CoSQA+, as evidenced by Cronbach’s Alpha of 0.956 and variance of 0.0004.
Our contributions can be summarized as follows:
  • We develop the CoSQA+ benchmark by pairing high-quality queries with high-quality multiple codes.
  • We enhance the construction process through candidate pair construction by multiple models, automated model annotation and code generation for unmatched queries.
  • We introduce the MMRR metric for assessing multi-code selection.
  • We publicly release CoSQA+ to facilitate further research in semantic code search.
Related Work
Code Search Datasets
Besides our work, there are other datasets and benchmarks in the realm of code search. Husain et al. (Husain et al. 2019) present CodeSearchNet which contains a wide range of query-code pairs from GitHub. CoSQA (Huang et al. 2021) and CS1QA (Lee, Seonwoo, and Oh 2022) contain the search logs collected from Microsoft Bing search engine but only cover Python and are limited in size. StaQC  (Yao et al. 2018) is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality. More recently, Li et al. (Li et al. 2024) introduce ProCQA, a large-scale community-based programming question answering dataset mined from StackOverflow with strict filtering strategies for quality and fairness control. This SOTA benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague.
Code Search Methods
The existing methods for code search can mainly be divided into two mainstreams: information retrieval (IR) based methods and deep learning (DL) based models. IR based methods extract keywords from queries and search for these keywords in code base to retrieve the most relevant code snippets (Bajracharya et al. 2006; Di Grazia and Pradel 2023; Zhang, Jin, and Zhou 2010). However, standard information retrieval methods have drawbacks in the code search domain, as they treat codes as text while programming languages and natural languages differ greatly. Fortunately, DL based methods receive great attention in NLP and achieve significant breakthroughs in code search research. The CodeSearchNet (Husain et al. 2019) dataset paves the way for CodeBERT (Feng et al. 2020) and GraphCodeBERT (Guo et al. 2020), both of which are bimodal pre-trained BERT (Devlin et al. 2018) based models that leverage Masked Language Modeling (MLM) and Replaced Token Detection (RTD) (Clark et al. 2020) objective. UniXcoder (Guo et al. 2022) unifies three pre-training designs into one architecture and utilizes AST structure and code comment to enhance the cross-modal alignment. Recent T5 (Raffel et al. 2020) based models like CodeT5 (Wang et al. 2021) and CodeT5+ (Wang et al. 2023) show outstanding performance in code search, utilizing the text-to-text transformer architecture and pre-training on large code datasets. While the two models achieve high performances in the CSN benchmark, GraphCodeBERT (Guo et al. 2020) demonstrates superior results by harnessing data flow encoded in the Abstract Syntax Tree (AST) of code to enrich the structural information during pre-training.
As large-scale models gain prominence, GPT (Radford et al. 2018) based models such as Codex (Chen et al. 2021), Claude 3 (Anthropic 2024) and GPT-4 (Achiam et al. 2023) are also employed in the field of code search, and exhibit remarkable performance.
Figure 1: The construction process of CoSQA+. The final CoSQA+ dataset includes queries from original CoSQA, and codes filtered from CodeSearchNet, StaQC and generated by Claude 3 Sonnet.
CoSQA+
Figure 1 illustrates the benchmark construction process for CoSQA+. The process starts with query and code collection, in which we leverage the queries from CoSQA and collect various code snippets from StaQC and CodeSearchNet. In candidate pairs construction, collected code snippets and queries are embedded as vectors to calculate similarity. For each query, the top 5 query-code pairs with the highest similarity are selected as candidates for potential matches. In model annotation, each query-code pair is annotated by Claude 3 Sonnet with “1” or “0”, indicating whether the code snippet exactly matches the query. These labeled pairs are then added to CoSQA+. In missing code generation, GPT-4o is used to generate codes for queries lacking exact matching code snippets. Each generated snippet is manually verified to ensure an exact match before incorporating the query-code pair into the CoSQA+. For manual verification, the annotators adhere strictly to the annotation guidelines to ensure consistency and accuracy.
Query and Code Collection
We take CoSQA as the query source, while CodeSearchNet Corpus (Husain et al. 2019) and StaQC (Yao et al. 2018) serve as the code sources. This section details the code data collection process, including the selection of appropriate code formats, sources, and a cleaning tool to promote the quality of the dataset.
Query Source
CoSQA queries are based on real user queries collected from Microsoft’s Bing search engine logs. We use these queries as the source for the following reasons. (1) We concentrate on code search in Python, and the filtered queries specifically target Python code search intent (Huang et al. 2021). (2) Utilizing real-world queries allows the code search model to be trained on data that reflects actual developer behavior. This helps the model effectively handle the diverse ways developers express their coding inquiries in real-world scenarios. (3) By using realistic queries, the code search model better meets user expectations, enhancing overall satisfaction. Developers are more likely to find relevant results when the queries they input closely align with their thought processes and communication about code (Li et al. 2020).
Code Format
The selection of code format is important for constructing a query-code matching dataset (Huang et al. 2021). Different formats include a statement (Yin et al. 2018), a code snippet or block (Yao et al. 2018), and a function (Husain et al. 2019). CoSQA simplifies the task and adopts a complete Python function as the answer, while we also adopt a code snippet or block for the following reasons. (1) Adopting a more inclusive code format allows for the collection of a broader dataset, which in turn can create a more comprehensive dataset. (2) This approach facilitates the evaluation of a model’s search capabilities across different code formats. (3) Diverse code examples play a crucial role in enhancing search accuracy by showcasing the variety of coding styles, syntax, and structural semantics  (Oliveira et al. 2023).
Code Source
To the best of our knowledge, CodeSearchNet and StaQC are the only origin datasets specifically designed for Python in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages (Husain et al. 2019). In CoSQA+, we keep all 1,156,085 Python functions. In addition, around 148K Python code snippets from StaQC are added to the candidate code pool of CoSQA+.
Code Cleaning
The 148K Python code base in StaQC contains many non-Python code snippets. Guesslang [^2] https://pypi.org/project/guesslang/ is a programming language detection tool based on Tensorflow. We utilize Guesslang to identify and filter out non-Python codes. The filtered-out codes can be categorized into four types: (1) incomplete snippets, (2) irrelevant fields, (3) codes written in other programming languages, and (4) others. More explanation and examples are shown in Appendix Figure 1. After cleaning, we retain 79k Python codes in StaQC.
Candidate Query-Code Pairs Construction
Labeling all query-code pairs is impractical. To improve efficiency, we only retain high-confidence pairs before annotation. This section details the construction process of high-confidence query-code pairs.
The relatedness between the query and code is quantified by the cosine similarity, calculated as:
cosine $\text{cosine similarity}=\frac{\mathbf{v}_{\text{query}}\cdot\mathbf{v}_{\text{code}}}{\|\mathbf{v}_{\text{query}}\|\|\mathbf{v}_{\text{code}}\|}$
where $\mathbf{v}_{\text{query}}$ and $\mathbf{v}_{\text{code}}$ are the embedding vectors of the query and code, respectively.
High similarity between query and code is fundamental for a high-quality dataset (Sun et al. 2022). During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. Specifically, we use three models: CodeBERT (Feng et al. 2020), UniXcoder (Guo et al. 2022) and CodeT5+ 110M embedding (Wang et al. 2023). By taking the average similarity score from these three models, we promote a more balanced and fair assessment of query-code pair quality.
Given our strict criteria for a complete match, only a few code snippets per query have the potential to be exact matches. To construct the candidate query-code pairs for further analysis, we select the top five code snippets with the highest average similarity scores for each query.
Model Annotation
Figure 2: Examples of query-code pairs.
The traditional data annotation process heavily relies on manual labor, demanding a deep understanding of Python and its libraries. Even experienced programmers need to use search tools to annotate query-code pairs, which is inefficient and time-consuming. In contrast, advanced language models like Claude 3 Sonnet (Anthropic 2024) offer an alternative. With comprehensive training and robust computational abilities, Claude 3 Sonnet can promote quality and accelerate the annotation process. Recognizing this potential, we implement a Chain of Thought (CoT) method to automate the annotation of each query-code pair. The CoT method integrates the specific guidelines directly into the model’s prompting mechanism. The performance of LLMs annotation is further evaluated in the experiment section.
Annotation Guidelines
During the annotation process, it is found that the majority of code snippets only partially fulfill the requirements of their corresponding queries. Some code snippets meet most of the query requirements but contain errors in certain details. Only a minority of snippets completely satisfy or even exceed all the requirements specified in the queries. To ensure fairness and consistency in our evaluations, it is crucial to maintain uniform judgment standards. Therefore, we categorize these situations into distinct cases to clarify the criteria and provide guidelines for annotators:
  1. 1.
    If the code can completely answer the query and even exceed the demand of the query, it is a correct answer. (Case (1), (2) in Figure  2)
  2. 2.
    If the code can satisfy most of the requirements of the query but unmeet a certain category of the query demand, it is a wrong answer. (Case (3) in Figure  2)
  3. 3.
    If the code only partially meets or even totally dismisses the demand of the query, it is a wrong answer. (Case (4) in Figure  2)
We compare our annotation guidelines with other datasets, such as CodeSearchNet (Husain et al. 2019) and CoSQA (Huang et al. 2021), and identify the key difference: these datasets deem codes correct if they meet just over half of the query’s demands. In contrast, our dataset adopts stricter labeling criteria, where only codes that fully satisfy the query’s requirements are labeled as correct. These stricter standards are implemented to better quantify the extent to which code meets the requirements of a query, thereby reducing inconsistencies in query-code pair annotations. Furthermore, stricter standards better align with the programmers’ requirements for retrieving more accurately matched query-code pairs in real-world scenarios.
Annotation
The task of manually annotating approximately 103K query-code pairs poses considerable logistical and financial challenges due to the extensive resources required for such an effort. To address this, we explore alternative strategies to reduce the burden without compromising the quality of annotations.
Our experiments reveal that the accuracy of labels generated by Claude 3 Sonnet closely approaches that of human annotators. Consequently, we opt to utilize Claude 3 Sonnet for automatic annotation. The prompts employed are based on a zero-shot chain of thought (CoT) approach, incorporating our annotation guidelines directly into the CoT input. This method ensures that the model not only understands the context of the task but also applies the specific criteria outlined in the guidelines when generating labels.
Figure 3: Prompt designed based on CoT.
Figure 3 illustrates how the guidelines are integrated with the CoT. This example provides a clear depiction of how the automated system processes the input. It produces a reliable annotation that meet the standards expected by human annotators. By leveraging the advanced prompting technique, we manage to maintain a near-human level of accuracy while significantly reducing the time and financial costs compared with manual annotation.
Missing Code Generation
After model annotation, we encountered a significant issue: 28.52% of the queries, amounting to 5,876, do not have exact matching code snippets. This presents a problem where high-quality queries are wasted since queries without exact matches cannot be used to calculate the MMRR or MRR. To address this issue and retain the high-quality queries within the CoSQA dataset for experimental purposes, we employ GPT-4o to generate code snippets that exactly match the remaining queries.
The generation process is followed by a manual inspection to ensure that the generated code meets the requirements of each query. If a generated code snippet does not fully satisfy the query’s needs, it is either regenerated or manually adjusted until it meets the criteria for an exact match.
This approach not only allows us to preserve the integrity and utility of our dataset but also ensures that all queries can contribute to the robustness of our experimental results. By leveraging the capabilities of GPT-4o, we effectively expand our dataset with high-quality, exact-match code snippets, thereby enhancing the overall quality and applicability of the CoSQA+ benchmark in real-world scenarios.
Benchmark Statistics
In total, 108,896 labeled query-code pairs are retained. The code snippets are from CSN, StaQC, and GPT-4o generation, while the queries are sourced from CoSQA. Table 1 shows the statistics of our CoSQA+ benchmark.
# of pairs # of code code avg. length
108,896 51,516 271.059
Table 1: Statistics of CoSQA+.
Evaluation
In this section, we conduct experiments with corresponding analysis to answer the following research questions (RQs):
RQ1: Is the quality of CoSQA+ higher than that of CoSQA?
RQ2: Can LLMs achieve similar results to humans in the annotation task?
RQ3: How effective are the existing approaches in semantic code search on our benchmark?
RQ4: Is MMRR reliable for evaluating models’ capacity in multi-choice code search?
RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
Experimental Setup
In the first experiment, we randomly select 1000 query-code pairs labeled “1” from CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, so we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
For the second experiment, we fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with MRR and MMRR as evaluation metrics. From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python. The queries in CSN99 are realistic queries sourced from Bing, carefully selected and verified by humans. In our training process, the model is trained using individual query-code pairs, similar to traditional one-to-one code search models. The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
Model Metric Fine-tune Dataset
None CoSQA CoSQA+
CodeBERT MRR 0.952 0.939 0.961
MMRR 0.882 0.844 0.899
UniXcoder MRR 0.891 0.951 0.954
MMRR 0.711 0.857 0.884
CodeT5+ MRR 0.794 0.938 0.953
embedding MMRR 0.504 0.860 0.889
Table 2: Performance of code search models fine-tuned with different data. The test set is CSN Python.
Results and Analysis
After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 62.9% of the better codes are from CoSQA+, while 37.1% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table 2, the three models fine-tuned with CoSQA+ outperforms those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in MMRR is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
RQ2: Annotation Performance of LLMs
To improve efficiency, Claude 3 Sonnet is used to annotate whether a query-code pair is an exact match in CoSQA+. This experiment aims to explore whether LLMs can achieve similar performance to humans in annotation.
Task
The annotation task is to determine whether a given query and the corresponding code snippet are an exact match, where LLMs need to classify each query-code pair as either an exact match (labeled “1”) or not (labeled “0”).
Metrics
This experiment can be divided into two parts. In the first part, we utilize Krippendorff’s alpha coefficients (Krippendorff 1980) $\alpha$ to measure the degree of agreement among LLM annotators and human annotators. For the second part, we use the metric Accuracy to evaluate the performance of humans and different LLMs.
Experimental Setup
As a test set, the Manual-CoSQA+ dataset consists of 1,000 manually labeled query-code pairs randomly selected from the original CoSQA+ dataset. According to the provided guidelines, each pair is manually labeled by three experienced programmers on exact match criteria and 50% match criteria respectively. Any disagreements are resolved by majority rule, setting the processed labels as the test standard.
For the models evaluation on Krippendorff’s alpha coefficients (Krippendorff 1980) $\alpha$ and accuracy, we select six LLMs: GPT-4o, GPT-4-turbo, GPT-3.5-turbo, Claude 3 Sonnet, Claude 3 Opus and Llama3-70B-instruct, as shown in Table 3 and Table 4. These models label the same pairs under identical guidelines.
Results and Analysis
Annotator 50% Match Exact Match
avg-human 0.86 0.89
GPT-4o 0.58 0.58
GPT-4-turbo 0.50 0.66
GPT-3.5-turbo 0.60 0.69
Claude 3 Sonnet 0.68 0.84
Claude 3 Opus 0.64 0.75
Llama3-70B-instruct 0.64 0.76
Table 3: Accuracy of different annotators on Manual-CoSQA+ dataset across match criteria. The 50% match and exact match criteria indicate that the code meets 50% or 100% of the query demands, respectively. Manual-CoSQA+ dataset includes 100 query-code pairs, each of which is annotated by 3 humans.
Model 50% Match Exact Match
GPT-4o 0.311 0.257
GPT-4-turbo 0.231 0.278
GPT-3.5-turbo 0.317 0.421
Claude 3 Sonnet 0.383 0.628
Claude 3 Opus 0.379 0.512
Llama3-70B-instruct 0.344 0.533
Table 4: Krippendorf’s Alpha Coefficients for different human-model combinations across match criteria.
The human annotator’s Krippendorff’s Alpha Coefficients are 0.354 and 0.607 for 50% match and exact match, respectively. As shown in Table 4, Claude 3 Sonnet outperforms the other models in this metric and is close to human levels. Therefore, it is selected to perform the three rounds of labeling the query-code pairs for the data annotation part. To guarantee the reliability and stability of labeling, Krippendorff’s alpha coefficients (Krippendorff 1980) are also calculated between different labeling rounds. Claude 3 Sonnet achieves a high score of 0.965 and 0.975, respectively. Other models show little difference in scores under the 50% match criteria, while varying significantly under the exact match criteria. This is because of the variations in their understanding of how codes can fully match the query.
As illustrated in Table 3, all models except GPT-4o have higher accuracy under the exact match criteria compared to the 50% criteria due to their similar performance on Krippendorff’s alpha with humans. Notably, GPT-4 Turbo and GPT-4o can catch semantic nuance between query and code, which may result in a far more strict judgment than humans in this annotation task. They focus intently on details, which misaligns with human tolerance for nuance and real-world code search demands.
RQ2 Summary: Using LLMs for annotation can achieve comparable results to manual annotation. The annotations by Claude 3 Sonnet are stable. Employing the exact matching criteria, as opposed to the 50% matching criteria, enhances annotation accuracy and consistency.
RQ3: Existing Methods Performance on CoSQA+
This experiment aims to evaluate the effectiveness of different models in accurately retrieving code snippets that match natural language queries on the CoSQA+ benchmark.
Task
The natural language code search task is a text retrieval problem. Given a natural language query $q_{i}$ and a collection of code snippets $C=\{c_{1},\ldots,c_{n}\}$, the goal is to identify all code snippets $\{c^{*}_{j}|c^{*}_{j}\in C\text{ and }c^{*}_{j}\text{ exactly matches }q_{i}\}$ that exactly match the query.
Metric
This task is assessed by the Mean Multi-choice Reciprocal Rank (MMRR), which can evaluate the capacity of different models in a relatively fair manner. Suppose there are $N$ queries and the correct number of code snippets corresponding to each query is k. We use $m_{i}$ to denote the average of the reciprocal ranks of $k$ terms for the $i$-th query. Then the MMRR can be defined as follows:
$MMRR=\frac{1}{N}\sum_{i=1}^{N}m_{i}$
where $m_{i}=\frac{1}{k}\cdot(\frac{1}{rank_{i1}}+\frac{1}{rank_{i2}-1}+...+\frac{1}{rank_{ik}-(k-1)})$. $rank_{ik}$ indicates the rank of the $k$-th correct code snippet in the model’s ranked list. For the $i$-th correct answer, we subtract $i-1$ from its rank to normalize the influence of multiple correct answers.
Experimental Setup
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark. The primary metric for evaluation is the MMRR, which is specifically designed for tasks where multiple correct solutions may exist for a given query. The MMRR better reflects the complexity and variability in code-related searches than the MRR does. The tested methods encompass both traditional information retrieval techniques like Lucene, Bag of Words(BoW) (Zhang, Jin, and Zhou 2010) and advanced models like CodeBERT (Feng et al. 2020), UniXcoder (Guo et al. 2022), CodeT5+ 110M embedding (Wang et al. 2023) and text-embedding-3-large [^3] https://openai.com/index/new-embedding-models-and-api-updates/.
Results and Analysis
Method MMRR MRR
Lucene 0.117 0.167
BoW 0.041 0.065
CodeBERT 0.288 0.392
UniXcoder 0.230 0.319
CodeT5+ embedding 0.181 0.266
Text-embedding-3-large 0.314 0.393
Table 5: Performance of Various Methods on CoSQA+ Using MMRR and MRR Metrics.
While traditional information retrieval methods are fast and less computationally intensive, Table 5 shows that they struggle with the semantic complexities required in accurately mapping queries to relevant codes in scenarios where multiple solutions are valid.
Advanced models like CodeBERT, UniXcoder, CodeT5+ and text-embedding-3-large significantly improve on MMRR and MRR, owing to training on a mix of natural and program languages. Text-embedding-3-large, as a text embedding model rather than one specifically designed for code search, achieves the best results among all the methods. This may be due to its higher dimensions and superior embedding quality. CodeBERT achieves results second only to Text-embedding-3-large. This performance is probably due to its encoder-only architecture and extensive pre-training on large pairs of natural and programming languages, especially Python, making it highly effective for code understanding.
RQ3 Summary: We present evaluation results of various methods on CoSQA+, demonstrating that deep learning-based models outperform traditional retrieval-based methods. Specifically, text-embedding-3-large achieves the highest performance among them.
RQ4: Is MMRR a Stable and Reliable Metric in Multi-Choice Code Search?
We conduct the following experiments to validate the stability and reliability of our proposed MMRR metric in evaluating models on the multi-choice code search task.
Metrics
We employ Cronbach’s Alpha (Tavakol and Dennick 2011) to measure the reliability of CoSQA+ in the MMRR. Additionally, we use average variance to evaluate the internal consistency and stability of the MMRR across different test sets and models. For stability evaluation, we calculate the average variance (Fisher 1936) of the MMRR across different models and test sets. The variance is computed by each model performance across all test sets, and then these variances are averaged to provide a single stability measure.
Experimental Setup
To evaluate the reliability and stability of MMRR in assessing models’ capability in multi-choice code search tasks, we exclude queries corresponding to only one correct code. The queries are divided into four distinct query bases, corresponding to the number of exactly matched codes (2, 3, 4, 5) for each query. The code base remains consistent with the original CoSQA+. This stratification allows us to explore how MMRR performs across queries of varying complexity and choice. Four different query bases form four test datasets with identical code base. We assess the multi-choice code search capabilities of three different models across these datasets.
Results and Analysis
MMRR achieves a high Cronbach’s Alpha of 0.956 and shows the lowest average variance at 0.0004, indicating high reliability and stability across all the test sets and models. These metrics confirm MMRR’s robustness in measuring multi-choice code search tasks under varying conditions.
RQ4 Summary: MMRR is a highly reliable and stable metric for evaluating models’ multi-choice code search capacities in CoSQA+.
Conclusion
In this paper, we present CoSQA+, a large-scale LLM-annotated query-code matching dataset, which contains 20,604 pairs of queries from original CoSQA, alongside diverse Python codes generated by LLMs and those filtered from StaQC and CSN. Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks. A novel metric, MMRR, is proposed to assess LLMs’ performance on multi-code selection. Experiments on the new benchmark show Claude 3 Sonnet excels in query-code pairs annotation tasks, while text-embedding-3-large is better suited for multi-choice code search tasks. We believe our LLM-annotated CoSQA+ dataset will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task.
Appendix A Codes filtered out from StaQC
Figure 4: Four types of codes filtered out from StaQC.
Guesslang is utilized to identify and remove non-Python codes in the StaQC dataset, as illustrated in Figure 4. The removed codes fall into four categories:
  • Incomplete snippets: These codes are in Python, but their semantics are not fully complete due to incomplete fragments.
  • Irrelevant fields: These snippets are not programming languages.
  • Other languages: These codes are not written in Python.
  • Others: These snippets are related to Python but include more than just codes.
Appendix B Missing code generation by GPT-4o
Figure 5: Examples of missing codes generated by GPT-4o
For queries lacking exact matches, we employ GPT-4o to generate potential code solutions, as illustrated in Figure 5
Appendix C Rationality of MMRR
Mean Multi-choice Reciprocal Rank metric is derived from MRR and extends the ranking of single choices to multiple choices for each query. In its definition, we adjust the rank with $i-1$ to normalize the impact of multiple relevant items, helping to even out the influence of each correct code snippet. Let’s consider the impact of multiple relevant codes on the reciprocal rank without adjustment.
Suppose there are two queries with different relevant code snippets. And A has 3 correct code snippets ranking 1, 2, and 3 in the list but another query B has 2 answers ranking 1, 2. We can compare the MMRR in different versions.
Without Adjustment:
For query A, the $m_{A}$ is $m_{A}=\frac{1}{3}\cdot(\frac{1}{rank_{A1}}+\frac{1}{rank_{A2}}+\frac{1}{rank_{A3}})=\frac{11}{18}$
For query B, $m_{B}$ can be $m_{B}=\frac{1}{2}\cdot(\frac{1}{rank_{B1}}+\frac{1}{rank_{B2}})=\frac{3}{4}$
$MMRR=\frac{1}{2}\cdot(\frac{11}{18}+\frac{3}{4})\approx 0.68$ At this point, the two values are different, but in reality, the model has correctly identified and ranked the correct codes for both queries at the top. Obviously, this calculation is unfair to query A.
With Adjustment:
For query A, the $m_{A}$ can be calculated as follows:$m_{A}=\frac{1}{3}\cdot(\frac{1}{rank_{A1}}+\frac{1}{rank_{A2}-1}+\frac{1}{rank_{A3}-2})=1$
For query B, its $m_{B}$ is $m_{B}=\frac{1}{2}\cdot(\frac{1}{rank_{B1}}+\frac{1}{rank_{B2}-1})=1$
$MMRR=\frac{1}{2}\cdot(1+1)=1$
In this case, $m_{A}$ and $m_{B}$ are equal and the MMRR increases, which is fair and reflects the model’s true capacity.
In summary, by incorporating the rank adjustment in the MMRR, we achieve a fairer and more reliable metric that accurately reflects the performance of models across queries with multiple correct answers.
Appendix D Performance Comparison of Various Models on CoSQA+ Using Different Metrics
Models MRR
2 3 4 5
UniXcoder 0.307 0.427 0.517 0.617
CodeT5+ 0.215 0.313 0.404 0.531
CodeBERT 0.405 0.497 0.571 0.651
Table 6: Performance Comparison of Various Models on CoSQA+ Using MRR.
Models MAP
2 3 4 5
UniXcoder 0.285 0.381 0.448 0.525
CodeT5+ 0.206 0.288 0.361 0.457
CodeBERT 0.374 0.416 0.493 0.554
Table 7: Performance Comparison of Various Models on CoSQA+ Using MAP.
Models MMRR
2 3 4 5
UniXcoder 0.192 0.213 0.229 0.273
CodeT5+ 0.124 0.141 0.159 0.193
CodeBERT 0.265 0.265 0.275 0.301
Table 8: Performance Comparison of Various Models on CoSQA+ Using MMRR.